xygen ( molecular weight 32) is compressed reversibly and polytropically in a


cylinder from 1.05 bar, 15oC to 4.2bar in such a way that one- third of the work input


is rejected as heat to the cylinder walls . calculate the final temperature of the oxygen.


Assume oxygen to be aperfect gas and take Cv = 0.649kJ/kg.K

1
Expert's answer
2022-04-06T13:59:56-0400

Answer

Temperature is given

"T'=T*(\frac{P'}{P}) ^(\frac{n-1}{n} ) \\=288*(\frac{420}{105}) ^(\frac{1.267-1}{1.267} ) \\=386K\\=113°C"